Aircraft arrive at the Mckinnon Airport with a Poisson distribution at a rate of one aircraft arrival every 25 minutes. What is the likelihood that at least 1 aircraft arrives in a 45-minute window of time?
Answer #1

The random variable X(t) = number of successes in time , follow Poisson distribution or X(t) ~ Poisson[lambda= at]

The pmf of X(t) is :

pk, N = __, x=0 1, 2, 3, 4 0,o/w

Given

X = Number of aircraft that arrive in t hours = 1 aircraft

a = 1 aircraft per 25 minute = 1 *60/25 aircraft per hour = 4 aircraft per hour

t = 45 minute = 0.75 hour

ous 0.75)

Now, probability that  at least 1 aircraft arrives in a 45-minute window of time:

P (X(t) 0.75) > 1)-1-PO)

eo, 75 × (0.75)0 0!

1e0.75

  =1 - 0.4734

  =0.5276